# Quantum resource degradation theory within the framework of observational entropy decomposition

### v2

@@ Abstract
We introduce a theory of quantum resource degradation grounded in a decomposition of observational entropy, which partitions the total resource into inter-block coherence ($\mathcal{C}_{\text{rel}}$) and intra-block noise ($\mathcal{D}_{\text{rel}}$). Under free operations, the total quantum resource is transformed into classical noise while its overall quantity remains conserved. We demonstrate that the metric $\eta$ functions as a diagnostic indicator, providing a new lens on optimization stagnation, particularly the barren plateau phenomenon (BPP) in variational quantum algorithms (VQAs). We substantiate this framework through rigorous mathematical analysis and numerical simulations, and we explore how these channels can be physically implemented in real quantum systems. Our approach offers a unified viewpoint on quantum thermalization, measurement-induced disturbance, and the degradation of quantum advantage in practical devices, while also improving optimization strategies for current and near-term noisy quantum hardware.
@@ I Introduction
Quantum technologies hold the potential for transformative progress by exploiting quantum resources such as coherence and entanglement. Quantum resource theory offers a rigorous mathematical framework to quantify these resources by distinguishing resource states from free states[Super, Coherence1, Resource, Coherence5, Measure1, Coherence6]. A foundational concept in this framework is monotonicity: resource measures must not increase under free operations. While this principle reliably describes the loss of resources, it also exposes a fundamental shortcoming: it does not reflect how the *quality* of those resources evolves.
This shortcoming has practical implications. In noisy quantum devices, algorithms frequently fail because coherence is effectively washed out by noise, yet conventional resource measures do not exhibit a corresponding sharp decline. This creates a paradox: we can quantify the amount of resource present, but not evaluate its quality. A quantum state may retain the same nominal resource “budget” even as its usefulness deteriorates, analogous to a currency that keeps its face value while its real purchasing power plummets.
We address this paradox by formulating a theory of quantum resource degradation grounded in a decomposition of observational entropy (OE). OE measures the uncertainty of a quantum state with respect to a specified coarse-grained measurement scheme[Dominik1, Dominik2, Dominik3, Correlation, Strasberg, Zhou5, Schindler2, Schindler3]. We demonstrate that this entropy can be structurally separated into two physically distinct components: inter-block coherence $\mathcal{C}_{\text{rel}}$ (a valuable quantum resource) and intra-block fluctuations $\mathcal{D}_{\text{rel}}$ (effectively classical noise). This decomposition reveals a generic degradation process: under free operations, $\mathcal{C}_{\text{rel}}$ is transformed into $\mathcal{D}_{\text{rel}}$, while their sum $\mathcal{O}_{\mathcal{C}}$ stays approximately invariant. In this way, quantum coherence is irreversibly converted into classical noise, maintaining the overall uncertainty but degrading its quantum nature.
@@
We put this theory to the test by examining a key unresolved problem in quantum computation: the barren plateau phenomenon (BPP) in variational quantum algorithms (VQAs)[VQA1, VQA2]. While earlier accounts have focused on entanglement or randomness, we instead pinpoint the loss of resource quality as the fundamental driver. Our resource-purity measure $\eta$ serves as an early-warning signal, forecasting when optimization is likely to fail before it actually does. By tracking $\eta$ during the algorithm’s execution, one can observe resource degradation in real time and take corrective measures before the variational optimization grinds to a halt.
@@
The structure of the paper is as follows. Section II reviews observational entropy and block coherence. In Section III, we present our decomposition framework together with the corresponding degradation mechanism. Section LABEL:Sec:5 contrasts our approach with existing methods. Section IV applies the framework to the analysis of barren plateaus in variational quantum algorithms. Finally, Section V explores broader implications and outlines possible directions for future work.
@@ II Background Knowledge
This section reviews the essential background needed to follow the theory developed in this work. Section II.1 introduces observational entropy (OE), and Section II.2 presents the framework of $\mathcal{C}_{\text{rel}}$.
@@ II.1 Observational Entropy
For a coarse-graining $\mathcal{C}=\{\hat{P}_{x}\}$ and an initial state $\rho$, the OE is defined as[Dominik1, Schindler2]
@@
Following Refs. [Zhou5, Schindler2, Schindler3], we define
@@
where $S(\rho)=-\operatorname{Tr}(\rho\log\rho)$ is the von Neumann entropy. This difference quantifies the state’s quantum coherence (or superposition) relative to the chosen coarse-graining. It satisfies[Zhou5]:
@@ II.2 Resource theory of inter-block coherence
The resource theory of block coherence, originally proposed by Åberg[Super], extends coherence theory from bases to general projective measurements. Let $\mathcal{H}=\bigoplus_{x}\mathcal{H}_{x}$ be a direct sum of orthogonal subspaces $\mathcal{H}_{x}$ with corresponding projectors $\hat{P}_{x}$. The set of block-incoherent (BI) states is
@@
$\Delta[\sigma]=\sum\limits_{x}\hat{P}_{x}\sigma\hat{P}_{x}.$ (4)
Maximally block-incoherent (MBI) operations $\Lambda_{\mathrm{MBI}}$ are quantum channels that map BI states to BI states[Iman1, Iman2, Iman3, Iman4, Iman5]:
@@ III.1 OE Decomposition
For a coarse-graining $\mathcal{C}$ and state $\rho$, we have the identity[Strasberg]
@@
Here $\rho_{x}=\hat{P}_{x}\rho\hat{P}_{x}/p_{x}$ and $\kappa_{x}=\hat{P}_{x}/V_{x}$ is the maximally mixed state in block $x$. The quantity $D(\rho\|\sigma)=\mathrm{Tr}[\rho(\log\rho-\log\sigma)]$ is the quantum relative entropy.
@@
where $\mathcal{D}_{\text{rel}}(\rho)=\sum_{x}p_{x}D[\rho_{x}\|\kappa_{x}]$.
@@ III.2 Resource Purity
The definition of $\eta$ is motivated by the one-way implication that if $\mathcal{O}_{\mathcal{C}}(\rho)=0$, then $\mathcal{C}_{\text{rel}}(\rho,\mathcal{C})=0$, while the converse does not hold[Zhou5]. The ratio $\eta$ quantifies what fraction of the observable effect $\mathcal{O}_{\mathcal{C}}$ is actually due to the useful resource. A larger value of $\eta$ therefore signals a higher-quality resource state, whereas a smaller value indicates that the resource is overwhelmed by intra-block noise. In this way, $\eta$ provides a quality-sensitive figure of merit for comparing resource states and for designing and optimizing distillation protocols across different physical platforms.
III.3 Nonconvexity of $\mathcal{O}_{\mathcal{C}}$
@@ III.4 Degradation Channels
We aim to demonstrate how resource quality degrades via coherence-to-purity conversion when the total resource is only approximately conserved. To this end, we construct explicit degradation channels. We support this mechanism with rigorous mathematical proofs and numerical examples, and we discuss how such channels can be physically realized in experimentally relevant quantum systems.
@@
In this work, we are particularly interested in regimes where, under a suitable class of free operations, coherence is irreversibly converted into purity, that is, $\mathcal{C}_{\text{rel}}$ decreases ($\Delta\mathcal{C}_{\text{rel}}<0$) while $\mathcal{D}_{\text{rel}}$ increases ($\Delta\mathcal{D}_{\text{rel}}>0$).
@@
[…]
@@
    • $\Delta[\rho]$: The block-dephasing map $\Delta$ removes all inter-block coherences and is a standard free operation.
[…]
@@
Proof: See Appendix AAppendix A: Proof of Proposition 1.
@@
  • Severe degradation ($\Delta\eta\geq 0.4$): The resource quality undergoes a strong deterioration, typically accompanied by a BPP.
@@
Proof: See Appendix BAppendix B: Proof of Theorem 1. We provide an analytic example using a 4-dimensional system and a specific channel $\Lambda_{\varepsilon}$. For $\varepsilon=0.26$, we obtain $\Delta\mathcal{C}_{\text{rel}}\approx-0.130$, $\Delta\mathcal{D}_{\text{rel}}\approx+0.129$, $\Delta\mathcal{O}_{\mathcal{C}}\approx-0.001$, and $\Delta\eta\approx 0.287$, which confirms the mechanism.
@@
Figure 1: Quantum resource measures as functions of the control parameter $\varepsilon$. In subplots (a) and (b), the blue curves represent $\mathcal{C}_{\text{rel}}$ and $\Delta\mathcal{C}_{\text{rel}}$, the red curves represent $\mathcal{D}_{\text{rel}}$ and $\Delta\mathcal{D}_{\text{rel}}$, the green curves represent $\mathcal{O}_{\mathcal{C}}$ and $\Delta\mathcal{O}_{\mathcal{C}}$, and the orange curve represents $\eta$. Panel (a) shows the behavior of the individual resource components as $\varepsilon$ is varied, while panel (b) shows the corresponding differential changes of the resource measures. Panel (c) quantifies the conversion rate between $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$. Panel (d) verifies the three conditions of Theorem 1 at the critical point. The control parameter is varied over $\varepsilon\in[0,0.5]$.
The evolution of the key physical quantities used in the proof of Theorem 1 (see Appendix Appendix B: Proof of Theorem 1) is shown in Fig. 1. When $\varepsilon$ approaches 1, the quantum state undergoes an almost complete collapse to $|00\rangle\langle 00|$. This extreme regime exhibits several features that lie outside the domain of interest for our theory.
Theorem 1 establishes the existence of degradation channels that implement the coherence-to-noise transformation. The proof does not rely on scanning the full parameter space; it suffices to demonstrate that such channels act as required within an appropriate parameter region. The essential degradation mechanism is already fully captured for $\varepsilon\in[0,0.26]$, so extending the analysis to larger $\varepsilon$ is not necessary to validate our theoretical claims. The observed conversion between coherence and purity in this interval provides direct evidence for the existence of the degradation channels described in Theorem 1. For clarity, we therefore restrict $\varepsilon$ to the interval $[0,0.5]$ in the main text.
Subplot (a) shows the evolution of the resource components as a function of $\varepsilon$. The quantity $\mathcal{O}_{\mathcal{C}}$ displays a convex dependence on $\varepsilon$ in $[0,0.26]$, while $\eta$ decreases monotonically. The convex behavior of $\mathcal{O}_{\mathcal{C}}$ signals a nonmonotonic degradation process: the resource does not decay linearly with the noise strength. The initial decrease corresponds to a temporary reduction in total inconsistency, while the subsequent recovery reflects compensating contributions from different resource components. Despite this variation, the value of $\mathcal{O}_{\mathcal{C}}$ at the end of the interval confirms the predicted approximate conservation of the total resource.
@@
Subplot (b) shows the changes in the individual resource measures. $\mathcal{C}_{\text{rel}}$ becomes negative, signaling decoherence. $\mathcal{D}_{\text{rel}}$ becomes positive, signaling the generation of purity. $\mathcal{O}_{\mathcal{C}}$ stays close to zero in the degradation region, demonstrating that the total resource is conserved during the transformation.
Subplot (c) quantifies the conversion rate between $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$. The efficiency $\Delta\mathcal{D}_{\text{rel}}/\lvert\mathcal{C}_{\text{rel}}\rvert$ approaches unity in certain regions, indicating nearly ideal conversion: each unit of lost quantum coherence is converted into one unit of gained classical purity. This behavior confirms the theoretical prediction that, under optimal conditions, quantum coherence is converted directly into classical noise.
Subplot (d) tests the three conditions of Theorem 1 at the critical point, showing that $\Delta\mathcal{C}_{\text{rel}}<0$, $\Delta\mathcal{D}_{\text{rel}}>0$, and $\Delta\mathcal{O}_{\mathcal{C}}\approx 0$ hold simultaneously. The pronounced drop in $\eta$ indicates a degradation in resource quality, while the near-constancy of $\mathcal{O}_{\mathcal{C}}$ confirms that the total amount of resource is conserved.

### v3

@@ Abstract
We introduce a theory of quantum resource degradation grounded in a decomposition of observational entropy, which partitions the total resource into inter-block coherence ($\mathcal{C}_{\text{rel}}$) and intra-block noise ($\mathcal{D}_{\text{rel}}$). Under free operations, the total quantum resource is transformed into classical noise while its overall quantity remains conserved. We demonstrate that the metric $\eta$ functions as a diagnostic indicator, providing a new lens on optimization stagnation, particularly the barren plateau phenomenon in variational quantum algorithms. We substantiate this framework through rigorous mathematical analysis and numerical simulations, and we explore how these channels can be physically implemented in real quantum systems. Our approach offers a unified viewpoint on quantum thermalization, measurement-induced disturbance, and the degradation of quantum advantage in practical devices, while also improving optimization strategies for current and near-term noisy quantum hardware.
@@ I Introduction
Quantum technologies hold the potential for transformative progress by exploiting quantum resources such as coherence and entanglement. Quantum resource theory offers a rigorous mathematical framework to quantify these resources by distinguishing resource states from free states[aberg2006quantifyingsuperposition, PhysRevLett.113.140401, PhysRevLett.123.110402, PhysRevA.102.012411, PhysRevA.104.012404, PhysRevA.93.032136]. A foundational concept in this framework is monotonicity: resource measures must not increase under free operations. While this principle reliably describes the loss of resources, it also exposes a fundamental shortcoming: it does not reflect how the quality of those resources evolves.
This shortcoming has practical implications. In noisy quantum devices, algorithms frequently fail because coherence is effectively washed out by noise, yet conventional resource measures do not exhibit a corresponding sharp decline. This creates a paradox: we can quantify the amount of resource present, but not evaluate its quality. A quantum state may retain the same nominal resource budget even as its usefulness gradually deteriorates, analogous to a currency that keeps its face value while its real purchasing power plummets.
We address this paradox by formulating a theory of quantum resource degradation grounded in a decomposition of observational entropy (OE). OE measures the uncertainty of a quantum state with respect to a specified coarse-grained measurement scheme[PhysRevA.99.010101, PhysRevA.99.012103, Safranek2021, PhysRevA.102.052407, PRXQuantum.2.030202, Zhou2023, Buscemi:2022wcd, Bai2024observational]. We demonstrate that this entropy can be structurally separated into two physically distinct components: inter-block coherence $\mathcal{C}_{\text{rel}}$ (a valuable quantum resource) and intra-block fluctuations $\mathcal{D}_{\text{rel}}$ (effectively classical noise). This decomposition reveals a generic degradation process: under free operations, $\mathcal{C}_{\text{rel}}$ is transformed into $\mathcal{D}_{\text{rel}}$, while their sum $\mathcal{O}_{\mathcal{C}}$ stays approximately invariant. In this way, quantum coherence is irreversibly converted into classical noise, maintaining the overall uncertainty but degrading its quantum nature.
@@
We put this theory to the test by examining a key unresolved problem in quantum computation: the barren plateau phenomenon (BPP) in variational quantum algorithms (VQAs)[McClean2018, Cerezo2021]. While earlier accounts have focused on entanglement or randomness, we instead pinpoint the loss of resource quality as the fundamental driver. Our resource-purity measure $\eta$ serves as an early-warning signal, forecasting when optimization is likely to fail before it actually does. By tracking $\eta$ during the algorithm’s execution, one can observe resource degradation in real time and take corrective measures before the variational optimization grinds to a halt.
@@
The structure of the paper is as follows. Section II reviews OE and block coherence. Section III, we present our decomposition framework together with the corresponding degradation mechanism. Section IV contrasts our approach with existing methods. Section V applies the framework to the analysis of barren plateaus in variational quantum algorithms. Finally, Section VI explores broader implications and outlines possible directions for future work.
@@ II Background Knowledge
This section reviews the essential background needed to follow the theory developed in this work.
@@ II.1 Observational Entropy
For a coarse-graining $\mathcal{C}=\{\hat{P}_{x}\}$ and an initial state $\rho$, the OE is defined as[PhysRevA.99.010101]
@@
Following Refs. [Zhou2023, Buscemi:2022wcd, Bai2024observational], we define
@@
where $S(\rho)=-\operatorname{Tr}(\rho\log_{2}\rho)$ is the von Neumann entropy. This difference quantifies the state’s quantum coherence (or superposition) relative to the chosen coarse-graining. It satisfies[Zhou2023]:
@@ II.2 Resource theory of inter-block coherence
The resource theory of block coherence, originally proposed by Åberg[aberg2006quantifyingsuperposition], extends coherence theory from bases to general projective measurements. Let $\mathcal{H}=\bigoplus_{x}\mathcal{H}_{x}$ be a direct sum of orthogonal subspaces $\mathcal{H}_{x}$ with corresponding projectors $\hat{P}_{x}$. The set of block-incoherent (BI) states is
@@
$\Delta(\sigma)=\sum\limits_{x}\hat{P}_{x}\sigma\hat{P}_{x}.$ (4)
Maximally block-incoherent (MBI) operations $\Lambda_{\mathrm{MBI}}$ are quantum channels that map BI states to BI states[Marvian_2013, Marvian2014, PhysRevA.80.012307, PhysRevA.93.052331, PhysRevA.94.052324]:
@@ III.1 OE Decomposition
For a coarse-graining $\mathcal{C}$ and state $\rho$, we have the identity[PRXQuantum.2.030202]
@@
Here $\rho_{x}=\hat{P}_{x}\rho\hat{P}_{x}/p_{x}$ and $\kappa_{x}=\hat{P}_{x}/V_{x}$ is the maximally mixed state in block $x$. The quantity $D(\rho\|\sigma)=\mathrm{Tr}[\rho(\log_{2}\rho-\log_{2}\sigma)]$ is the quantum relative entropy.
@@
where $\mathcal{D}_{\text{rel}}(\rho)=\sum_{x}p_{x}D[\rho_{x}\|\kappa_{x}]$. In this work, we refer to $\mathcal{O}_{\mathcal{C}}(\rho)$ as the total inconsistency.
@@ III.2 Resource Purity
The definition of $\eta$ is motivated by the one-way implication that if $\mathcal{O}_{\mathcal{C}}(\rho)=0$, then $\mathcal{C}_{\text{rel}}(\rho,\mathcal{C})=0$, while the converse does not hold[Zhou2023]. The ratio $\eta$ quantifies what fraction of the observable effect $\mathcal{O}_{\mathcal{C}}$ is actually due to the useful resource. A larger value of $\eta$ therefore signals a higher-quality resource state, whereas a smaller value indicates that the resource is overwhelmed by intra-block noise. In this way, $\eta$ provides a quality-sensitive figure of merit for comparing resource states and for designing and optimizing distillation protocols across different physical platforms.
III.3 Nonconvexity of Total Inconsistency
@@ III.4 Degradation Channels
We aim to demonstrate how resource quality degrades via coherence-to-noise conversion, even when the total resource is only approximately conserved. To this end, we construct explicit degradation channels. We support this mechanism with rigorous mathematical proofs and numerical examples, and we discuss how such channels can be physically realized in experimentally relevant quantum systems.
@@
In this work, we are particularly interested in regimes where, under a suitable class of free operations, coherence is irreversibly converted into noise, that is, $\mathcal{C}_{\text{rel}}$ decreases ($\Delta\mathcal{C}_{\text{rel}}<0$) while $\mathcal{D}_{\text{rel}}$ increases ($\Delta\mathcal{D}_{\text{rel}}>0$).
@@
where $\mathcal{D}_{\beta}(\rho)=\beta\Delta(\rho)+(1-\beta)\sum\limits_{x}\hat{P}_{x}|\psi_{0}\rangle\langle\psi_{0}|\hat{P}_{x}$. The parameter $\alpha$ controls the overall noise strength of the channel, while $\beta$ tunes the relative weight of the two noise mechanisms, thus specifying the pathway by which quantum coherence is converted into classical noise.
@@
    • $\Delta(\rho)$: The block-dephasing map $\Delta$ removes all inter-block coherences and is a standard free operation.
    • $\sum_{x}\hat{P}_{x}|\psi_{0}\rangle\langle\psi_{0}|\hat{P}_{x}$: This map prepares a block-diagonal state by projecting the target ground state $|\psi_{0}\rangle$ into the different blocks. Such block-diagonal state preparation is free[Marvian_2013].
@@
Proof: See Appendix VII.
@@
  • Severe degradation ($\Delta\eta\geq 0.4$): The resource quality undergoes a strong deterioration.
@@
Proof: See Appendix VIII. We provide an analytic example using a 4-dimensional system and a specific channel $\Lambda_{\varepsilon}$. For $\varepsilon=0.26$, we obtain $\Delta\mathcal{C}_{\text{rel}}\approx-0.130$, $\Delta\mathcal{D}_{\text{rel}}\approx+0.129$, $\Delta\mathcal{O}_{\mathcal{C}}\approx-0.001$, and $\Delta\eta\approx 0.287$, which confirms the mechanism.
@@
Figure 1: In Fig. 1(a) and Fig. 1(b), the blue curves represent $\mathcal{C}_{\text{rel}}$ and $\Delta\mathcal{C}_{\text{rel}}$, the red curves represent $\mathcal{D}_{\text{rel}}$ and $\Delta\mathcal{D}_{\text{rel}}$, the green curves represent $\mathcal{O}_{\mathcal{C}}$ and $\Delta\mathcal{O}_{\mathcal{C}}$, and the orange curve represents $\eta$. Fig. 1(a) shows the behavior of the individual resource components as $\varepsilon$ is varied, while Fig. 1(b) shows the corresponding differential changes of the resource measures. Fig. 1(c) quantifies the conversion rate between $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$. Fig. 1(d) verifies the three conditions of Theorem 1 at the critical point. The control parameter is varied over $\varepsilon\in[0,0.5]$.
The evolution of the key physical quantities used in the proof of Theorem 1 (see Appendix VIII) is shown in Fig. 1. When $\varepsilon$ approaches 1, the quantum state undergoes an almost complete collapse to $|00\rangle\langle 00|$. This extreme regime exhibits several features that lie outside the domain of interest for our theory.
Theorem 1 establishes the existence of degradation channels that implement the coherence-to-noise transformation. The proof does not rely on scanning the full parameter space. It suffices to demonstrate that such channels act as required within an appropriate parameter region. The essential degradation mechanism is already fully captured for $\varepsilon\in[0,0.26]$, so extending the analysis to larger $\varepsilon$ is not necessary to validate our theoretical claims. The observed conversion between coherence and noise in this interval provides direct evidence for the existence of the degradation channels described in Theorem 1. For clarity, we therefore restrict $\varepsilon$ to the interval $[0,0.5]$ in the main text.
Fig. 1(a) shows the evolution of the resource components as a function of $\varepsilon$. The quantity $\mathcal{O}_{\mathcal{C}}$ displays a convex dependence on $\varepsilon$ in $[0,0.26]$, while $\eta$ decreases monotonically. The convex behavior of $\mathcal{O}_{\mathcal{C}}$ signals a non-monotonic degradation process: the resource does not decay linearly with the noise strength. The initial decrease corresponds to a temporary reduction in total inconsistency, while the subsequent recovery reflects compensating contributions from different resource components. Despite this variation, the value of $\mathcal{O}_{\mathcal{C}}$ at the end of the interval confirms the predicted approximate conservation of the total resource.
@@
Fig. 1(b) shows the changes in the individual resource measures. $\mathcal{C}_{\text{rel}}$ becomes negative, signaling decoherence. $\mathcal{D}_{\text{rel}}$ becomes positive, signaling the generation of purity. $\mathcal{O}_{\mathcal{C}}$ stays close to zero in the degradation region, demonstrating that the total resource is conserved during the transformation.
Fig. 1(c) quantifies the conversion rate between $\mathcal{C}_{\text{rel}}$ and $\mathcal{D}_{\text{rel}}$. The efficiency $\Delta\mathcal{D}_{\text{rel}}/\lvert\mathcal{C}_{\text{rel}}\rvert$ approaches unity in certain regions, indicating nearly ideal conversion: each unit of lost quantum coherence is converted into one unit of gained classical noise. This behavior confirms the theoretical prediction that, under optimal conditions, quantum coherence is converted directly into classical noise.
Fig. 1(d) tests the three conditions of Theorem 1 at the critical point, showing that $\Delta\mathcal{C}_{\text{rel}}<0$, $\Delta\mathcal{D}_{\text{rel}}>0$, and $\Delta\mathcal{O}_{\mathcal{C}}\approx 0$ hold simultaneously. The pronounced drop in $\eta$ indicates a degradation in resource quality, while the near-constancy of $\mathcal{O}_{\mathcal{C}}$ confirms that the total amount of resource is conserved.